# Mitigating Error Accumulation in Co-Speech Motion Generation via Global Rotation Diffusion and Multi-Level Constraints

### v1

@@ Abstract
Reliable co-speech motion generation requires precise motion representation and consistent structural priors across all joints. Existing generative methods typically operate on local joint rotations, which are defined hierarchically based on the skeleton structure. This leads to cumulative errors during generation, manifesting as unstable and implausible motions at end-effectors. In this work, we propose GlobalDiff, a diffusion-based framework that operates directly in the space of global joint rotations for the first time, fundamentally decoupling each joint’s prediction from upstream dependencies and alleviating hierarchical error accumulation. To compensate for the absence of structural priors in global rotation space, we introduce a multi-level constraint scheme. Specifically, a joint structure constraint introduces virtual anchor points around each joint to better capture fine-grained orientation. A skeleton structure constraint enforces angular consistency across bones to maintain structural integrity. A temporal structure constraint utilizes a multi-scale variational encoder to align the generated motion with ground-truth temporal patterns. These constraints jointly regularize the global diffusion process and reinforce structural awareness. Extensive evaluations on standard co-speech benchmarks show that GlobalDiff generates smooth and accurate motions, improving the performance by 46.0% compared to the current SOTA under multiple speaker identities. Project page: https://xiangyue-zhang.github.io/GlobalDiff
@@ Introduction
At the motion level, we propose to use temporal embeddings extracted via a multi-scale variational encoder (Kingma, Welling et al. 2013; Ma et al. 2025) to represent the motion-level structure. By aligning the temporal features of the generated and ground-truth sequences, we ensure dynamic consistency, rhythm synchronization, and smooth temporal transitions throughout the motion.
@@
Diffusion-based Motion Generation. Diffusion models have become a dominant framework in human motion generation, particularly in text-to-motion and co-speech motion tasks, due to their strong generative capacity. For co-speech motion, DiffuseStyleGesture (Yang et al. 2023) pioneered expressive audio-conditioned gesture synthesis, incorporating rhythm alignment and style control. DiffSHEG (Chen et al. 2024b) extended these approaches by jointly modeling facial and body gestures efficiently. GestureDiffuCLIP (Ao, Zhang, and Liu 2023) further introduced multimodal style control using CLIP embeddings. HoloGest (Cheng and Huang 2025) leveraged motion priors and decoupled body parts during diffusion. LivelySpeaker (Zhi et al. 2023) proposed a two-stage method, generating semantic-aware gestures from text before rhythm refinement via audio-driven diffusion. RAG-Gesture (Mughal et al. 2025) introduced retrieval-augmented generation for explicit semantic alignment. However, previous methods overlook the error accumulation problem caused by local rotations and the forward kinematics process, resulting in significant inconsistencies in the motion of end-effectors. In this work,we predict global joint rotations directly and introduce multi-level constraints—at the joint, skeleton, and motion levels—to ensure structural consistency and expressive realism.
@@
Datasets. For training and evaluation, we adopt the BEAT2 dataset (Liu et al. 2024a), which provides approximately 60 hours of 3D motion data paired with speech from 25 speakers (12 female, 13 male). It contains 1,762 dialogue sequences, each lasting around 65 seconds on average. While prior methods typically report results on a fixed subset of speakers, we evaluate our model on both the full test set and the subset corresponding to speaker ”Scott” to support fair comparison with single-speaker setups and assess generalization across diverse speakers, as recommended by (Mughal et al. 2025).
@@
Implementation Details. Our model is trained on four NVIDIA V100 GPUs for 1,000 epochs with a batch size of 128, taking approximately 17 hours. We use the ADAM optimizer with a learning rate of 1e-4. We empirically set the number of virtual nodes to 6. Following (Liu et al. 2024a), we train the model starting with an 8-frame seed pose. For streaming, the last 8 frames from the previous clip are reused as the seed for the next, enabling long-form generation with only the initial 8 frames required.
Metrics. To evaluate the quality of generated gestures, we use several metrics targeting different aspects of performance. For overall realism, we adopt the Fréchet Gesture Distance (FGD) (Yoon et al. 2020), which compares the statistical distribution of generated gestures to that of the ground truth (GT). Motion diversity (Diversity) is measured using (Li et al. 2021a), defined as the mean L1 distance between pairs of generated clips. Temporal coordination between speech and motion is assessed with the Beat Alignment (BeatAlign) metric (Li et al. 2021b), which reflects how well gesture timing aligns with the rhythm of the accompanying audio. For facial accuracy, we report vertex MSE (Yang et al. 2023), measuring deviations between predicted and ground-truth facial vertices.
@@
Qualitative Comparisons. As shown in Figure 3, GlobalDiff consistently generates semantically grounded and physically coherent co-speech motions across all speaker identities, outperforming baselines. We compare gestures produced by all three methods on four speakers—Wayne (ID1), Scott (ID2), Lawrence (ID4), and Carla (ID6)—across varied utterances. RAG-GESTURE (Mughal et al. 2025) is excluded due to unavailable public outputs. For the phrase “exaggerating”, GlobalDiff produces wide, symmetric arm extensions that clearly express emphasis, whereas EMAGE shows constrained gestures and HoloGest yields flat, non-expressive arm poses. For “wow”, our model generates elevated open-hand motions that reflect surprise, while EMAGE collapses the upper limbs and HoloGest shows minimal variation. When expressing “driving”, GlobalDiff synthesizes realistic, steering-like motions. In contrast, EMAGE and HoloGest often fall back on generic, low-effort gestures. For “all the time”, our model maintains consistent chest-level hand motions across speakers, preserving semantic clarity, while EMAGE introduces lateral imbalance and HoloGest shows reduced motion fidelity. This comparison underscores the advantage of our global rotation strategy and structural constraints in producing expressive, semantically accurate, and identity-consistent gestures.
@@ 1 Speaker All Speakers
Table 2: Ablation study on the effectiveness of each component within the GlobalDiff on the speaker 2.
@@
Figure 7: Results of the user study. GlobalDiff received higher preference across all evaluation metrics, with particularly strong scores in realism and semantic consistency.
@@ Conclusion
Acknowledgments. This work was supported by Alibaba Research Intern Program. The numerical calculation is supported by Tongyi Lab, Alibaba Group.

### v2

@@ Abstract
Reliable co-speech motion generation requires precise motion representation and consistent structural priors across all joints. Existing generative methods typically operate on local joint rotations, which are defined hierarchically based on the skeleton structure. This leads to cumulative errors during generation, manifesting as unstable and implausible motions at end-effectors. In this work, we propose GlobalDiff, a diffusion-based framework that operates directly in the space of global joint rotations for the first time, fundamentally decoupling each joint’s prediction from upstream dependencies and alleviating hierarchical error accumulation. To compensate for the absence of structural priors in global rotation space, we introduce a multi-level constraint scheme. Specifically, a joint structure constraint introduces virtual anchor points around each joint to better capture fine-grained orientation. A skeleton structure constraint enforces angular consistency across bones to maintain structural integrity. A temporal structure constraint utilizes a multi-scale variational encoder to align the generated motion with ground-truth temporal patterns. These constraints jointly regularize the global diffusion process and reinforce structural awareness. Extensive evaluations on standard co-speech benchmarks show that GlobalDiff generates smooth and accurate motions, improving the performance by 46.0% compared to the current SOTA under multiple speaker identities.
Projecthttps://xiangyue-zhang.github.io/GlobalDiff
@@ Introduction
At the motion level, we propose to use temporal embeddings extracted via a multi-scale variational encoder (Kingma, Welling et al. 2013; Ma et al. 2025; Xiao et al. 2025; Yu et al. 2025) to represent the motion-level structure. By aligning the temporal features of the generated and ground-truth sequences, we ensure dynamic consistency, rhythm synchronization, and smooth temporal transitions throughout the motion.
@@
Diffusion-based Motion Generation. Diffusion models are widely used for human motion generation, including text-to-motion and co-speech tasks. DiffuseStyleGesture (Yang et al. 2023) first demonstrated expressive audio-conditioned gestures with rhythm and style control. Later works extended this direction: DiffSHEG (Chen et al. 2024b) modeled facial and body motions jointly; GestureDiffuCLIP (Ao, Zhang, and Liu 2023) introduced CLIP-based style control; HoloGest (Cheng and Huang 2025) incorporated motion priors with part-wise diffusion; LivelySpeaker (Zhi et al. 2023) used a two-stage semantic-then-audio pipeline; and RAG-Gesture (Mughal et al. 2025) added retrieval for stronger semantic alignment. Yet these methods rely on local joint rotations, where forward kinematics causes error growth along the kinematic chain. This leads to unstable end-effectors. We address this by predicting global joint rotations directly and applying joint-, skeleton-, and motion-level constraints to maintain structural consistency and expressive quality.
@@
Datasets. For training and evaluation, we adopt the BEAT2 dataset (Liu et al. 2024a), which provides approximately 60 hours of 3D motion data. It contains 1,762 dialogue sequences, each lasting around 65 seconds on average. While prior methods typically report results on a fixed subset of speakers, we evaluate our model on both the full test set and the subset corresponding to speaker ”Scott” to support fair comparison with single-speaker setups and assess generalization across diverse speakers, as recommended by (Mughal et al. 2025).
@@
Implementation Details. Training is performed on four NVIDIA V100 GPUs for 1,000 epochs with a batch size of 128 and takes about 17 hours. We use the ADAM optimizer with a learning rate of 1e-4. The number of virtual nodes is set to 6. Following (Liu et al. 2024a), training begins with an 8-frame seed pose. For streaming, the last 8 frames of each clip serve as the seed for the next, so long sequences require only the initial 8 frames.
Metrics. We evaluate motion using four criteria. Fréchet Gesture Distance (FGD) (Yoon et al. 2020) measures distribution-level realism. Diversity (Li et al. 2021a) captures variation across generated clips through mean L1 differences. Beat Alignment (BeatAlign) (Li et al. 2021b) assesses rhythm synchrony between speech and motion. Facial accuracy is measured with vertex MSE (Yang et al. 2023).
@@
Qualitative Comparisons. Figure 3 shows that GlobalDiff produces semantically clear and physically stable co-speech motions across all speaker identities. We compare our model with EMAGE and HoloGest on four speakers—Wayne (ID1), Scott (ID2), Lawrence (ID4), and Carla (ID6). RAG-GESTURE (Mughal et al. 2025) is omitted due to unavailable public results. For phrases such as “exaggerating” and “wow”, GlobalDiff generates wide, well-formed arm and hand motions that reflect emphasis and surprise, whereas EMAGE tends to restrict movement and HoloGest often produces flat poses. For “driving”, our model synthesizes steering-like motions, while the baselines revert to generic gestures. For “all the time”, GlobalDiff maintains stable chest-level motions across speakers, while EMAGE introduces imbalance and HoloGest reduces motion detail. These observations highlight the value of global rotation prediction and our structural constraints.
@@ 1 Speaker All Speakers
Table 2: Ablation study on the effectiveness of each component in GlobalDiff on speaker 2
@@
Figure 7: Results of the user study.
@@ Conclusion
Acknowledgments
This work was supported by Alibaba Research Intern Program. The numerical computation was supported by Tongyi Lab, Alibaba Group